OpenStudy (anonymous):

Evaluate the improper integral e^x with limits 0 and -ve infinity.

OpenStudy (anonymous):

\[\int\limits_{-\Pi}^{0}e^x dx\]

OpenStudy (anonymous):

1

OpenStudy (dumbcow):

well the anti-derivative of e^x is e^x then evaluate from -inf to 0 \[= e^{0} - e^{-\infty} = 1 - 0 = 1\]

OpenStudy (dumbcow):

\[e^{-\infty} = \frac{1}{e^{\infty}} = \frac{1}{\infty} = 0\]

OpenStudy (anonymous):

Thanks!!

OpenStudy (dumbcow):

no problem